how many litters of a 40% alcohol solution must be mixed with 65% solution to obtain 20 liters of a 50% solution?
Answer by mananth(16946)

---------------percent ---------------- quantity
Alcohol I 40 ---------------- x liters
Alcohol II 65 ---------------- 20 - x liters
Mixture 50.00% ---------------- 20

40x+65( 20 - x ) = 50 * 20

40x + 1300 - 65 x = 1000
40x - 65 x = 1000 - 1300
-25 x = -300
/ -25
x = 12 liters 40.00% Alcohol I
8 liters 65.00% Alcohol II